OpenStudy (anonymous):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {14, 18, 21, 15, 17} Data Set 2: {15, 17, 22, 20, 16}

OpenStudy (anonymous):

I think its 1.5

OpenStudy (anonymous):

@iGreen

OpenStudy (youngstudier):

Ok, so what is the mean for each data set?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (youngstudier):

I am literally doing this stuff right now in my math course.

OpenStudy (anonymous):

umm xD

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

for set 1: 17 and Set 2: 18

OpenStudy (youngstudier):

Ok, gimme one second, I have to look up the formula in my book and I'll double check the things as we go through them.

OpenStudy (anonymous):

ok xD

OpenStudy (anonymous):

Wait the answer would be 1.1?

OpenStudy (anonymous):

Because you divide 18 by 17

OpenStudy (youngstudier):

urm, one moment, I'm helping two other people as well so I gotta check the means quick

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I know its that because i retook them Because i didnt pass it the first time Btw not a test But i did get to retake it

OpenStudy (youngstudier):

Ok, so the means are right, now we have to find the MAD

OpenStudy (anonymous):

But the questions did change

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

how do you get mad?

OpenStudy (anonymous):

For first one it is 2 and 2nd one it is 2.4

OpenStudy (youngstudier):

I'll double check it, one moment

OpenStudy (anonymous):

MAD - Mean absolute deviation of a component of a data set is the total difference between that component and a given point. In general, the central tendency is the measure of a point from which the deviation is measured, more frequently the median or sometimes the mean of the data set. The Mean Absolute Deviation also called as average deviation of a data set {x1, x2, ..., xn} is the average of the absolute deviations and is a abstract statistic of statistical distribution or set of data. The mean absolute deviation often abbreviated as MAD and calculated from the formula

OpenStudy (anonymous):

Total Inputs(N) =(15, 17, 22, 20, 16) Total Inputs(N)=5 Mean(xm)= (x1+x2+x3...xN)/N Mean(xm)= 90/5 Means(xm)= 18 ------------------------------------------- MAD= 1/(N)*(|x1-xm|+|x2-xm|+..+|xN-xm|) =1/5(|15-18|+|17-18|+|22-18|+|20-18|+|16-18|) =1/5(3+1+4+2+2)) =2.4

OpenStudy (youngstudier):

Ok, so your MAD are correct

OpenStudy (anonymous):

:)

OpenStudy (youngstudier):

now we subtract the first data set mean from the second data set mean

OpenStudy (youngstudier):

so 17-18=?

OpenStudy (youngstudier):

@MasterTeemo

OpenStudy (anonymous):

Ok back @YoungStudier

OpenStudy (anonymous):

-1

OpenStudy (youngstudier):

ok, so now we divide that by the larger MAD

OpenStudy (anonymous):

6.4?

OpenStudy (anonymous):

so 6.4/-1?

OpenStudy (youngstudier):

mm, no. Hold on.

OpenStudy (anonymous):

wait it was 2.4 xD

OpenStudy (youngstudier):

So it would be -1/2.4=?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

-0.41666666666

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

0.4? is the answer?

OpenStudy (youngstudier):

Do you have any answer choices for this one?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its a type in

OpenStudy (youngstudier):

does it say to round to the nearest number or somethin?

OpenStudy (youngstudier):

@MasterTeemo ?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

expressed as a decimal to the nearest tenth?

OpenStudy (youngstudier):

One moment

OpenStudy (anonymous):

would it have been 18-17 then 1/2.4?

OpenStudy (anonymous):

then the answer be 0.4?

OpenStudy (anonymous):

or no?

OpenStudy (anonymous):

http://ncalculators.com/statistics/mean-absolute-deviation-calculator.htm is what i used to get the MAD

OpenStudy (youngstudier):

Nope, -0.42 www.calculatorsoup.com/calculators/math/roundingnumbers.php type the number into the calculator

OpenStudy (anonymous):

But i has to be -0.4 xD

OpenStudy (anonymous):

it*

OpenStudy (youngstudier):

did you type it in?

OpenStudy (anonymous):

yep

OpenStudy (youngstudier):

so it is -0.42

OpenStudy (anonymous):

-0.4

OpenStudy (anonymous):

Because the bar thing has a dropdown choice you you have to pick tenths

OpenStudy (youngstudier):

Oh, sorry, I didn't realize. Yes, -0.4 is correct

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

XD only on question 2 Now time for question 3 :)

OpenStudy (youngstudier):

LOL, I forgot about the option. m'kay, do you need help with that one also?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

3/5 XD

OpenStudy (anonymous):

Alexander found the means-to-MAD ratio of two data sets to be 0.8. What can he conclude about the distributions? A. They are similar. B. They are somewhat similar. C. They are different. D. They are identical.

OpenStudy (anonymous):

ill open new questions and tag you and give you 5 medals k?

OpenStudy (youngstudier):

ok...

OpenStudy (anonymous):

xD dont worry i wont lie like most people do

OpenStudy (youngstudier):

Oh no, I've never had that happen to me. I don't actually really care that much about the medals LOL

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

So The answer would be B for this next q?

OpenStudy (youngstudier):

Ok, so what do you think the answer is to the second one you posted here?

OpenStudy (anonymous):

B?

OpenStudy (youngstudier):

mmm, no.

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

oh wait its 0.8 so no

OpenStudy (anonymous):

would it be Identical?

OpenStudy (youngstudier):

Remember: If the ratio is less than or equal to one, the two distributions are similar. What do you think the answer is now?

OpenStudy (anonymous):

D?

OpenStudy (youngstudier):

I don't remember reading anything about ratios being identical in the lesson, do you?

OpenStudy (anonymous):

nope xD because i really didnt read it

OpenStudy (anonymous):

i kinda read part but not all

OpenStudy (anonymous):

so C? xD

OpenStudy (youngstudier):

Reread what I said a little bit above. It's not C

OpenStudy (anonymous):

oh ok So A?

OpenStudy (anonymous):

xD sorry wasnt paying attenchen

OpenStudy (youngstudier):

Yep! A is correct. XD I'm no iGreen, but hey, I actually read through this lesson (I haven't done the bookwork or test yet. LOL)

OpenStudy (anonymous):

Some of my friends just said it was D xD

OpenStudy (anonymous):

majesty777: Nope D, they are identical because it says TWO sets of data of 0.8

OpenStudy (youngstudier):

Hm..... Okay....Now you made me confused. O_O

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

majesty777: Well it says two sets of the SAME 0.8 that is identical

OpenStudy (anonymous):

Who DO I Beleve?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

ok ill go with a